# Light Scalars in Light of UV/IR Mixing Classicalization via Synergy between Vainshtein & Chameleon Screenings

### v1

@@ 2.1 Classicalization in a Nutshell
Like BHs undergoing Hawking radiation, the classicalon evaporates thermally into $N_{\circledast}\gg 1$ quanta.
@@ 2.2.2 Classicalon Solution from Classical Vainshtein Screening
Here, $J$ represents a pointlike source [^12] In dark energy theories, the magnitude $\mathcal{E}$ of the localized source is represented by the mass of the astrophysical object (e.g., a star) around which Vainshtein screening takes place [3]. Naturally, the pointlike approximation should be relaxed when required. of magnitude $\mathcal{E}>0$. The underlying concept is that if a classicalon can be generated by a pointlike external source, it will necessarily also arise from the kinetic self-interaction when attempting to sharply localize a wave packet of $\phi$-bosons. It is important to note that, while the source term $\mathcal{L}_{J}$ preserves shift symmetry, it explicitly breaks Lorentz-Poincaré invariance and the $\mathbb{Z}_{2}$ symmetry.
@@
The background solution $\overline{\phi}^{\,\prime}(r)$ in these regimes is expressed as
[…]
where a real solution in the non-linear regime exists only if $c_{2}=-1$. Notably, for all regimes,
@@ 2.2.3 Radiative Stability from Quantum Vainshtein Screening
    For a given value of $\mathcal{E}$, the factor $Z_{\phi}(r_{0})$ increases as $r_{0}\to 0$, strengthening the blueshift effect as one approaches the source. Notably, $\overline{\phi}^{\,\prime}(r)$ diverges for $r_{0}\to 0$, reflecting the pointlike nature of the external source $J(x)$.
@@ 2.2.4 Classicalization from Kinetic Self-interactions
$\forall\,r>R_{\circledast}\,,\quad\phi(r)\sim-\frac{1}{\Omega r}\cdot\frac{\mathcal{E}_{0}}{\Lambda_{\ast}},$ (2.34)
@@
$S_{X}=\int d^{4}x\ \mathcal{L}_{X},\quad\mathcal{L}_{X}=\frac{1}{2}\,\partial^{\mu}\phi\partial_{\mu}\phi+\frac{G_{\phi}}{2}\left(\partial^{\mu}\phi\partial_{\mu}\phi\right)^{2}.$ (2.37)
@@ 2.3.3 Causality and Positivity Bounds: The Fate of the Fate
Classicalization embodies the notion of a minimal length scale that an interaction can probe and is thus non-perturbatively non-local, despite the apparent locality of the k-essence Lagrangian density in Eq. (2.9). This can be formalized within the axiomatic framework of QFT using Jaffe’s classification of strictly localizable, quasi-local, and non-localizable fields, specifically by examining the ultraviolet behavior of the spectral density functions [192, 193, 194, 134, 195]. The Wightman formulation, causality, and the standard properties of the $S$-matrix for theories with non-localizable fields have been studied in Refs. [196, 197, 198, 199, 200, 201, 202, 203, 204, 205, 206, 207, 208, 209, 210]. Positivity bounds are typically derived from dispersion relations for textbook QFTs, which deal with tempered localizable fields, a subclass of strictly localizable fields [195], whereas classicalizing QFTs involve non-localizable fields [134, 135]. The status of positivity bounds on QFTs with non-localizable fields is discussed in Refs. [211, 195, 135, 212], and the results for classicalizing theories are consistent with the violation of the bounds obtained by assuming tempered localizable fields [195, 135], e.g., $c_{2}=-1$. The properties of non-localizable fields, whose spectral density functions grow faster than those of strictly localizable fields, explain why classicalization appears at odds with any interpretation in a Wilsonian perspective [148]. We also stress that this is the reason the classicalization proposal for scalars is fundamentally different from the controversial Higgsplosion scenario [213, 214, 215, 216, 217] in the framework of the standard electroweak theory, involving only strictly localizable fields [218, 219], while both deal with multi-scalar boson production at high energies.
@@ 2.3.4 Non-renormalizable Operators Demystified
This EFT is equivalent to the original theory only at tree level [^18] More generally, a systematic method for integrating out fields in perturbation theory is the background field method [220]. for the $\Phi$-field, as the non-local form factor in the quartic term encodes the tree-level propagation of the massive $\Phi$-particle. For Euclidean momenta $p_{E}^{2}\ll M^{2}$, a Taylor expansion of this form factor—known as the operator product expansion (OPE)—produces a local EFT Lagrangian density:
@@
where the tower of effective operators is truncated at the appropriate order in $\square/M^{2}$ via power counting. These non-renormalizable operators [^19] While this terminology is standard, it is not accurate, as the EFT remains perturbatively renormalizable at any given order in perturbation theory. thus represent the low-energy remnants of the $\Phi$-propagator, with the interaction scale $\Lambda$ set by the heavy mass $M$ if $y_{\psi}\sim 1$.
@@
As a final remark on the interpretation of UV-screeners within the EFT framework, the reader may question how the values of the couplings $\kappa_{i}$ align with the renormalization program. A detailed analysis of the perturbative renormalization of the EFT governing the fluctuations around the classicalon background lies beyond the scope of this article (see Ref. [153] for an attempt in this direction). Here, we offer only preliminary considerations on the subject. Within the classicalon radius, the running of these couplings is effectively frozen, as the dynamics in this region is predominantly classical, as demonstrated in Ref. [151]. Since quantum corrections are suppressed, no large logarithms should require resummation. Consequently, one should be able to choose the renormalized values of the couplings $\kappa_{i}$ to match those of the classical Lagrangian $(\hbar\to 0)$, with, e.g., a subtraction point specified by the background amplitude $\overline{\phi_{0}}^{\prime}\equiv\Lambda_{\ast}^{2}$. The counterterms should then be employed to absorb the UV divergences, as usual in the background field method [1]. This also implies that if an operator vanishes classically, it is incorporated with a renormalized coupling that is zero in the Lagrangian density, yet accompanied by a non-vanishing counterterm to address the UV divergences [153].
@@ 3.1 Massive K-essence
Here, we define the renormalized field $\delta\phi_{Z}(x)$ from Eq. (2.28) to identify the effective mass $\overline{m}(r_{0})$ of the fluctuations of the background $\overline{\phi}(r)$. Deep within the Vainshtein core, where $Z_{\phi}(r_{0})\gg 1$, the mass undergoes a redshift, yielding $\overline{m}(r_{0})\ll m$. Consequently, the mass term does not disturb the screening mechanism, and the background solution is perturbatively stable.
@@ 3.2.3 Radiative Stability and Classicalization
which must remain stable [^23] It is important to note, in general, that fluctuations of the background may exhibit tachyonic instabilities. Instead of ghosts, this is not pathological [3]: classicalon solutions do not correspond to stable field configurations; instead, they decay into a large number of soft quanta. under quantum fluctuations of the background, $\delta\phi(x)=\phi(x)-\overline{\phi}(r)$. To demonstrate this stability, we observe that this asymptotic form can be expressed as
@@
From this discussion, it is evident [^24] For $|\overline{\phi}|\gg\Lambda_{\phi}$, the exponential factor in Eq. (3.22) always dominates over both polynomial and logarithmic factors of $|\overline{\phi}|/\Lambda_{\phi}$. that radiative corrections cannot alter the form of the potential for $|\phi|\gg\Lambda_{\phi}$ in a manner that would induce vacuum instability (see Ref. [234] for a review on false vacuum decay). This chameleon screening can therefore serve as a stabilization mechanism for the vacuum in theories involving scalar fields [^25] For discussions on vacuum stability in another class of non-local theories, see Refs. [235, 236]..
@@ 3.3.1 Yukawa Coupling vs Vainshtein Screening
By selecting a potential $V(\phi)$ compatible with Vainshtein screening, the background solution $\overline{\phi}(r)$ for $r\ll R_{C}$ is provided by Eq. (2.16), and we add the perturbations $\delta\phi(x)=\phi(x)-\overline{\phi}(r)$. The fermionic Lagrangian density is then modified to
$\mathcal{L}_{\Psi}=\frac{i}{2}\,\overline{\Psi}\gamma^{\mu}\overleftrightarrow{\partial_{\mu}}\Psi-\overline{M_{\Psi}}(r)\,\delta\phi_{Z}\overline{\Psi}\Psi-\overline{y}(r)\,\overline{\Psi}\Psi,$ (3.26)
@@
for a moderately weak Yukawa coupling, and $R_{V}\gg r\gg\ell_{\ast}$. The fluctuations consequently receive substantial quantum corrections from the Yukawa coupling, which destabilizes the classicalon solution. The issue becomes even more pronounced within the chameleon halo, where $Z_{\phi}\sim 1$. Thus, the model requires modification.
@@ 3.3.2 Conformal Coupling and Chameleon Screening
Consequently, a background with $|\phi|\gg\Lambda_{\Psi}$ leads to a suppressed coupling $y_{C}[\phi]$, reflecting another manifestation of the underlying chameleon screening mechanism. Like $V(\phi)$, this operator serves as another example of a term that breaks the shift symmetry—otherwise preserved by the UV-screener responsible for classicalization—and hinders Vainshtein screening, thereby requiring suppression via chameleon screening.
@@ 4 Conclusion and Outlook
This framework offers a fresh perspective on the existence of light scalar bosons, when they appear to suffer from a little hierarchy problem: the UV/IR mixing intrinsic to the classicalization phenomenon requires such a hierarchy to facilitate a consistent self-UV-completion of the theory. Nevertheless, it is important to emphasize that the specific choices of UV-screeners—including the scalar potentials and conformal couplings—serve merely as illustrative examples of the underlying screening mechanisms. Exploring alternative possibilities in this direction therefore remains a valuable endeavor.

### v2

@@ 2.1 Classicalization in a Nutshell
Like BHs undergoing Hawking radiation, the classicalon evaporates thermally into $N_{\circledast}\gg 1$ quanta (see Fig. 3).
@@ 2.2.2 Classicalon Solution from Classical Vainshtein Screening
Here, $J$ represents a pointlike source [^12] In dark energy theories, the magnitude $\mathcal{E}$ of the localized source is represented by the product of a scalar coupling and the mass of the astrophysical object (e.g., a star) around which Vainshtein screening takes place [3]. Naturally, the pointlike approximation should be relaxed when required. of magnitude $\mathcal{E}>0$. The underlying concept is that if a classicalon can be generated by a pointlike external source, it will necessarily also arise from the kinetic self-interaction when attempting to sharply localize a wave packet of $\phi$-bosons. It is important to note that, while the source term $\mathcal{L}_{J}$ preserves shift symmetry, it explicitly breaks Lorentz-Poincaré invariance and the $\mathbb{Z}_{2}$ symmetry.
@@
To ensure a real solution that smoothly interpolates between the linear and non-linear regimes, the background solution $\overline{\phi}^{\prime}(r)$ takes the form
$\dfrac{\overline{\phi}^{\prime}(r)}{\Lambda_{\ast}^{2}}=\left\{\begin{array}[]{cllll}\left(\dfrac{R_{V}}{r}\right)^{2}&\ll 1&\text{for}&r\gg R_{V}&\text{(linear regime)},\\[5.0pt] \mathcal{O}(1)&&\text{for}&r\sim R_{V}&\text{(transition regime)},\\[5.0pt] (-c_{2})^{1/3}\left(\dfrac{R_{V}}{r}\right)^{2/3}&\gg 1&\text{for}&r\ll R_{V}&\text{(non-linear regime)},\end{array}\right$. (2.16)
which admits a real solution only when $c_{2}=-1$. Notably, for all regimes,
@@ 2.2.3 Radiative Stability from Quantum Vainshtein Screening
    For a given value of $\mathcal{E}$, the factor $Z_{\phi}(r_{0})$ increases as $r_{0}\to 0$, strengthening the blueshift effect as one approaches the source. Notably, $\overline{\phi}^{\,\prime}(r)$ diverges for $r_{0}\to 0$, reflecting the pointlike nature of the external source $J(x)$. Moreover, for the purposes of the present discussion, we need not assume the validity of our semiclassical treatment—or, consequently, of the background solution in Eq. (2.16)—for $r\lesssim\ell_{\ast}$.
@@ 2.2.4 Classicalization from Kinetic Self-interactions
$\forall\,r\gg R_{\circledast}\,,\quad\phi(r)\sim-\frac{1}{\Omega r}\cdot\frac{\mathcal{E}_{0}}{\Lambda_{\ast}},$ (2.34)
@@
$S_{X}=\int d^{4}x\ \mathcal{L}_{X},\quad\mathcal{L}_{X}=\frac{1}{2}\,\partial^{\mu}\phi\partial_{\mu}\phi+\frac{G_{\phi}}{4}\left(\partial^{\mu}\phi\partial_{\mu}\phi\right)^{2}.$ (2.37)
@@ 2.3.3 Causality and Positivity Bounds: The Fate of the Fate
Classicalization embodies the notion of a minimal length scale that an interaction can probe and is thus non-perturbatively non-local, despite the apparent locality of the k-essence Lagrangian density in Eq. (2.9). This can be formalized within the axiomatic framework of QFT using Jaffe’s classification of strictly localizable, quasi-local, and non-localizable fields, specifically by examining the UV behavior of the spectral density functions [192, 193, 194, 134, 195]. The Wightman formulation, causality, and the standard properties of the $S$-matrix for theories with non-localizable fields have been studied in Refs. [196, 197, 198, 199, 200, 201, 202, 203, 204, 205, 206, 207, 208, 209, 210]. Positivity bounds are typically derived from dispersion relations for textbook QFTs, which deal with tempered localizable fields, a subclass of strictly localizable fields [195], whereas classicalizing QFTs involve non-localizable fields [134, 135]. The status of positivity bounds on QFTs with non-localizable fields is discussed in Refs. [211, 195, 135, 212], and the results for classicalizing theories are consistent with the violation of the bounds obtained by assuming tempered localizable fields [195, 135], e.g., $c_{2}=-1$. The properties of non-localizable fields, whose spectral density functions grow faster than those of strictly localizable fields, explain why classicalization appears at odds with any interpretation in a Wilsonian perspective [148]. We also stress that this is the reason the classicalization proposal for scalars is fundamentally different from the controversial Higgsplosion scenario [213, 214, 215, 216, 217] in the framework of the standard electroweak theory, involving only strictly localizable fields [218, 219], while both deal with multi-scalar boson production at high energies.
@@ 2.3.4 Non-renormalizable Operators Demystified
This EFT is equivalent to the original theory only at tree level [^18] More generally, a systematic method for integrating out fields in perturbation theory is the background field method [220]. for the $\Phi$-field, as the non-local form factor in the quartic term encodes the tree-level propagation of the massive $\Phi$-particle. For Euclidean momenta $p_{E}^{2}\ll M_{\Phi}^{2}$, a Taylor expansion of this form factor—known as the operator product expansion (OPE)—produces a local EFT Lagrangian density:
@@
where the tower of effective operators is truncated at the appropriate order in $\square/M^{2}$ via power counting. These non-renormalizable operators [^19] While this terminology is standard, it is not accurate, as the EFT remains perturbatively renormalizable at any given order in perturbation theory. thus represent the low-energy remnants of the $\Phi$-propagator, with the interaction scale $\Lambda$ set by the heavy mass $M_{\Phi}$ if $y_{\psi}\sim 1$.
@@
As a final remark on the interpretation of UV-screeners within the EFT framework, the reader may question how the values of the couplings $\kappa_{i}$ align with the renormalization program. A detailed analysis of the perturbative renormalization of the EFT governing the fluctuations around the classicalon background lies beyond the scope of this article (see Ref. [153] for an attempt in this direction). Here, we offer only preliminary considerations on the subject. Within the classicalon radius, the running of these couplings is effectively frozen, as the dynamics in this region is predominantly classical, as demonstrated in Ref. [151]. Since quantum corrections are suppressed [^21] In certain cases, non-renormalization theorems guarantee that a coupling $\kappa_{i}$ does not run [225, 226]., no large logarithms should require resummation. Consequently, one should be able to choose the renormalized values of the couplings $\kappa_{i}$ to match those of the classical Lagrangian $(\hbar\to 0)$, with, e.g., a subtraction point specified by the background amplitude $\overline{\phi_{0}}^{\prime}\equiv\Lambda_{\ast}^{2}$. The counterterms should then be employed to absorb the UV divergences, as usual in the background field method [1]. This also implies that if an operator vanishes classically, it is incorporated with a renormalized coupling that is zero in the Lagrangian density, yet accompanied by a non-vanishing counterterm to address the UV divergences [153].
@@ 3.1 Massive K-essence
Here, we use the renormalized field $\delta\phi_{Z}(x)$ from Eq. (2.28) to identify the effective mass $\overline{m}(r_{0})$ of the fluctuations of the background $\overline{\phi}(r)$. Deep within the Vainshtein core, where $Z_{\phi}(r_{0})\gg 1$, the mass undergoes a redshift, yielding $\overline{m}(r_{0})\ll m$. Consequently, the mass term does not disturb the screening mechanism, and the background solution is perturbatively stable.
@@ 3.2.3 Radiative Stability and Classicalization
which must remain stable under quantum fluctuations of the background, $\delta\phi(x)=\phi(x)-\overline{\phi}(r)$. To demonstrate this stability, we observe that this asymptotic form can be expressed as
@@
From this discussion, it is evident [^24] For $|\overline{\phi}|\gg\Lambda_{\phi}$, the exponential factor in Eq. (3.22) always dominates over both polynomial and logarithmic factors of $|\overline{\phi}|/\Lambda_{\phi}$. that radiative corrections cannot alter the form of the potential for $|\phi|\gg\Lambda_{\phi}$ in a manner that would induce vacuum instability [^25] It is important to note, in general, that fluctuations of the background may exhibit tachyonic instabilities. Instead of ghosts, this is not pathological [3]: classicalon solutions do not correspond to stable field configurations; instead, they decay into a large number of soft quanta. (see Ref. [236] for a review on false vacuum decay). The combination of Vainshtein and chameleon screenings can therefore serve as a stabilization mechanism for the vacuum in theories involving scalar fields [^26] For discussions on vacuum stability in another class of non-local theories, see Refs. [237, 238]..
@@ 3.3.1 Yukawa Coupling vs Vainshtein Screening
By selecting a potential $V(\phi)$ compatible with Vainshtein screening, the background solution $\overline{\phi}(r)$ for $r\ll R_{C}$ is provided by Eq. (2.16), and we add the perturbations $\delta\phi(x)=\phi(x)-\overline{\phi}(r)$. The fermionic Lagrangian density is then written as
$\mathcal{L}_{\Psi}=\frac{i}{2}\,\overline{\Psi}\gamma^{\mu}\overleftrightarrow{\partial_{\mu}}\Psi-\overline{M_{\Psi}}(r)\,\overline{\Psi}\Psi-\overline{y}(r)\,\delta\phi_{Z}\overline{\Psi}\Psi,$ (3.26)
@@
for a moderately weak Yukawa coupling, and $R_{V}\gg r\gg\ell_{\ast}$. The fluctuation mass consequently receives substantial quantum corrections from the Yukawa coupling, which destabilizes the classicalon solution. The issue becomes even more pronounced within the chameleon halo, where $Z_{\phi}\sim 1$. Thus, the model requires modification.
@@ 3.3.2 Conformal Coupling and Chameleon Screening
Consequently, a background with $|\phi|\gg\Lambda_{\Psi}$ leads to a suppressed coupling $y_{C}[\phi]$, reflecting another manifestation of the underlying chameleon screening mechanism. Like $V(\phi)$, the Yukawa interaction serves as another example of a term that breaks the shift symmetry—otherwise preserved by the UV-screener responsible for classicalization—and hinders Vainshtein screening, thereby requiring suppression via chameleon screening.
@@ 4 Conclusion and Outlook
This framework offers a fresh perspective on the existence of light scalar bosons, when they appear to suffer from a little hierarchy problem: the UV/IR mixing intrinsic to the classicalization phenomenon requires such a hierarchy to facilitate a consistent self-UV-completion of the theory. Nevertheless, it is important to emphasize that our specific choices of UV-screeners—including the scalar potentials and conformal couplings—serve merely as illustrative examples of the underlying screening mechanisms. Exploring alternative possibilities in this direction therefore remains a valuable endeavor.